OpenStudy (anonymous):

I'll Give Out A Medal! Please Someone HELP! Solve. c minus 9 is greater than twelve c is greater than 3 c is greater than 9 c is greater than twelve c is greater than twenty-one

OpenStudy (anonymous):

If you can solve c - 9 = 12, then you can solve c - 9 > 12. Just replace the "=" with ">"

OpenStudy (anonymous):

So Its Either B Or C?

OpenStudy (anonymous):

can you solve c - 9 = 12 ?

OpenStudy (anonymous):

Yeah Give Me A Sec.

OpenStudy (anonymous):

Its 21 - 9 = 12

OpenStudy (anonymous):

@BangkokGarrett

OpenStudy (anonymous):

Someone Help! :c

OpenStudy (ivettef365):

you have c - 9 > 12 just add 9 on both sides what do you get ?

OpenStudy (anonymous):

your answer is c = 12 to be precise. that's correct. now replace the = with a >. There's your answer.

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

So Its C?

OpenStudy (anonymous):

add 9 to both sides c - 9 + 9 > 12 + 9

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good luck. gotta go

OpenStudy (ivettef365):

so what is c -9 +9 > 12 + 9 ???